OpenStudy (anonymous):

I need some help with this question. Use the properties of real numbers to rewrite the expression. I'll put the expression in the comments.

OpenStudy (anonymous):

\[(2\5) * (-7) * (5/2)\]

OpenStudy (mathstudent55):

A number times its reciprocal equals 1. For example, 1/2 * 2/1 = 1 1/3 * 3/1 = 1 2/3 * 3/2 = 1

OpenStudy (mathstudent55):

Also, the commutative property of multiplication allows you to change the order of the factors. Examples: 2 * 5 = 5 * 2 7 * 3 = 3 * 7

OpenStudy (mathstudent55):

The associative property of multiplication allows you to group factors in different ways. Examples: (2 * 3) * 5 = 2 * (3 * 5) 3 * (x * y) = (3 * x) * y

OpenStudy (anonymous):

So would the answer be -7?

OpenStudy (mathstudent55):

Now look at your fractions. They are 2/5 and 5/2. Group them together and multiply them together. Then multiply the result by -7.

OpenStudy (mathstudent55):

Correct, since 2/5 * 5/2 = 1 since 2/5 and 5/2 are reciprocals. 1 times any number is the number, so (2/5) * (-7) * (5/2) = = (2/5) * (5/2) * (-7) = 1 * (-7) = -7

OpenStudy (mathstudent55):

Good job!

OpenStudy (anonymous):

Thank you! I got it right. I got an 80% because I missed one of my questions on my assignment. But I got this one right! Thank you!

OpenStudy (mathstudent55):

You're welcome.